OpenStudy (anonymous):

Find the 3x3 matix X that satisfies the matrix equation XA = B, where A=| 1 1 1 | and B = | 1 3 0 | |-1 0 -2| |-2 1 -2 | |1 0 -1| |1 4 -1|

OpenStudy (anonymous):

[3 2 -1] [1 2 -1] [4 2 1]

OpenStudy (anonymous):

how would you solve it?

OpenStudy (anonymous):

do u know how to do matrix multiplication?

OpenStudy (anonymous):

yes, I just don't know how would u find X.

OpenStudy (anonymous):

so u see: 1(x)+1(y)+1(z)=1 1(x)+0(y)+0(z)=3 1(x)-2(y)-1(z)=0 solving x,y and z gives 3, 2 and -1 respectively which is the first row of the matrix X

OpenStudy (anonymous):

I see, I figured it out. I did the transpose of A augmented with the transpose of B and got an answer and transpose the answer to get the solution.